OpenStudy (anonymous):

find the measure of each angle of the triangle (6x+10) (x+15) (x-5)

OpenStudy (kc_kennylau):

They add up to \(180^\circ\)!

OpenStudy (anonymous):

\[6x+10+x+15+x-5=180\]

OpenStudy (anonymous):

\[8x-20=180\]

OpenStudy (anonymous):

\[8x=180-20\]

OpenStudy (anonymous):

\[8x=160\]

OpenStudy (anonymous):

\[x=20\]

OpenStudy (anonymous):